1) Which substance will heat up faster:

Aluminum (s) specific heat = 0.90 J/g°C

Brass (s) specific heat = 0.380 J/g°C

2)A student mixes 50.0 mL of a 0.100 M AgNO3 solution with 50.0 mL of a 0.100 M HCl solution in a coffee cup calorimeter. Upon mixing the temperature of the resulting solution increases from 22.30°C to 23.11°C. If we assume the resulting solution has a final volume of 100.0 mL and the density and specific heat of the resulting solution are the same as that of water, what is the ΔH of the reaction in kJ/mol AgNO3 reacted?

AgNO3 + HCl → AgCl + HNO3
Answer #1

Question 1

Aluminum (s) specific heat = 0.90 J/g°C

Brass (s) specific heat = 0.380 J/g°C

The metal which has lower specific heat capacity will heat up faster. Hence Brass will heat up faster.

Question 2

we can use the following formula

Q = mc∆T

Q = heat energy (Joules, J), m = mass of a substance (g)

c = specific heat (units J/g∙°C), is a symbol meaning "the change in"

∆T = change in temperature (°C Celcius)

Q = ? m = 100 ml = 100 g C = 4.184 J/g∙°C  ∆T = 23.11 °C - 22.3 °C = 0.81 °C

Q = 100 g x 4.184 J/g∙°C x 0.81 °C

Q = 338.904 Joules = 0.3388 kJ

Let us assume there is no heat absorbed by calorimeter

Moles of AgNO3 used = 50 x 0.1 / 1000 = 0.005 Moles

ΔH of the reaction =  0.3388 kJ /0.005 Moles =   67.78 kJ /mol

Hence ΔH of the reaction is 67.78 kJ /mol